OpenStudy (anonymous):

I'm stuck right here: (x^2-(sqrt(x^2+4))(2+sqrt(x^2+4)))/(4x(sqrt(x^2+4))(2+sqrt(x^2+4)))

hartnn (hartnn):

\(\Large \dfrac{(x^2-(\sqrt{(x^2+4)})(2+\sqrt{(x^2+4)}))}{(4x(\sqrt{(x^2+4)})(2+\sqrt{(x^2+4)}))}\) i can see \((2+\sqrt{(x^2+4)})\) getting cancelled...

hartnn (hartnn):

or did you purposely multiplied numerator and denominator by \(2+\sqrt{x^2+4}\) ?

OpenStudy (anonymous):

The one that should be multiplied in the numerator is (sqrt(x^2+4) and (2+sqrt(x^2+4))

hartnn (hartnn):

may i know the original question ? :)

OpenStudy (anonymous):

Here:

hartnn (hartnn):

so it'll be \(\Large \dfrac{x^2-(\sqrt{(x^2+4)})(2+\sqrt{(x^2+4)})}{(4x(\sqrt{(x^2+4)})(2+\sqrt{(x^2+4)}))}\) lets try to simplify the numerator \(x^2-(\sqrt{(x^2+4)})(2+\sqrt{(x^2+4)}) = x^2 - 2\sqrt{(x^2+4)} -\sqrt{(x^2+4)}\sqrt{(x^2+4)} \)

hartnn (hartnn):

= \(x^2 -2\sqrt{(x^2+4)} - (x^2 +4)\) notice the x^2 term gets cancelled.... makes sense till now?

OpenStudy (anonymous):

Yes

hartnn (hartnn):

so what does numerator simplify to ?

OpenStudy (anonymous):

it will become -2(sqrt(x^2+4)) -4

hartnn (hartnn):

correct and what if you factor out -2 from that ?

OpenStudy (anonymous):

-2(sqrt(x^2+4) +2)

hartnn (hartnn):

so we now have \(\large \dfrac{-2((2+\sqrt{(x^2+4)})}{(4x(\sqrt{(x^2+4)})(2+\sqrt{(x^2+4)}}\) and now \((2+\sqrt{(x^2+4)}\) gets cancelled! :)

OpenStudy (anonymous):

Thanks you!!!! :D

hartnn (hartnn):

welcome ^_^